User name Markowitz
Log entry time 15:52:40 on July19,2003
Entry number 104927
keyword=day shift summary
08:00 started shift with N2 reference cell running, 90 psig, run 2467, 22467, 3357.
08:20 set fields for HRSR to 2.005 and septum magnet to 185.50 magnet current readback (185.75 set current). Also doing EPR/NMR measurements.
08:45 Finished EPR/NMR, moving to N2 reference cell (140 psig). Magnets at 2.005 GeV/c.
08:58 Start runs 22468/2468/3358.
09:08 Move to no target.
09:16 Ask for raster off, do harp scan 2167. Get horizontal sigma of 213 um and vertical of 115 um.
09:23 raster on, empty target 22469/2469/3359
09:40 raster off, empty target 22470/2470/3360
shower counters seem badly gain matched in hardware and possibly two channels overflow the ADC? I expanded the histogram ADC range to 8000. The histograms
are NOT corrected properly, and the hardware voltages are badly mismatched but the spectra does not overflow the ADC. This means the detector information could be recovered in software analysis.
09:55 move to 3He polarized
Runs 22470-4/2470-2475/3360-3364 are production, 3He polarized 180 degrees, beam half-wave plate out.
11:45 ask MCC to insert beam half-wave plate
11:47 Start runs 22475/2475/3365. Production 3He polarized 180 degrees, beam half-wave plate in.
Runs 22475-80, 2475-2480, 3365-70 all 3He polarized at 180 degrees, beam half-wave plate in.
13:32 change to reference cell and set target polarization to 0.
13:44 Run 22481/2481/3371 Empty reference cell with 10 uA current, raster on.
13:56 Change PS: T1=5, T2=5, left arm to T3=2 (from T3=1) and start run 22482/2482/3372
14:10 RUNS 22483/2483/3373 empty reference with 2 uA, raster on, PS1=1, PS2=1, PS3=1
14:25 change to 3He polarized, 0 degrees, half-wave plate in
14:21 Change PS1,PS2 back to 10 each
14:31 start runs 22484/2484/3374 3He polarized, raster on, 2 uA, half-wave plate in
runs 22484-22488, 2484-2488, and 3374-3378 are all 3He polarized at 0 degrees, half-wave plate in.